OpenStudy (anonymous):

What is the positive solution for 5^2-9x=0 ?

OpenStudy (saifoo.khan):

is it 5x^2 - 9x = 0?

OpenStudy (anonymous):

It's 5x squared - 9x = 0

OpenStudy (saifoo.khan):

its answer is 9/5 & 0.

OpenStudy (anonymous):

But there should only be one answer...the positive one. So, what is it?

OpenStudy (saifoo.khan):

no, there will b 2 answers.

OpenStudy (anonymous):

I got 25/9

OpenStudy (saifoo.khan):

no, its 9/5 & 0.

OpenStudy (anonymous):

Yeah. Your right Saifoo but then the positive one should be it....there should be a negative anser and a positive one.

OpenStudy (anonymous):

*answer

OpenStudy (saifoo.khan):

no, its not imp.

OpenStudy (anonymous):

???

OpenStudy (saifoo.khan):

there will the just theses two answers.

OpenStudy (saifoo.khan):

these*

OpenStudy (anonymous):

Is it 5^2 or 5x^2 in the origianl problem?

OpenStudy (anonymous):

okay...lol. here's whats up.

OpenStudy (saifoo.khan):

it is 5x^2

OpenStudy (saifoo.khan):

i think u typed the wrong ques

OpenStudy (anonymous):

i have a project due tommorrow for math I and i need someone to help tutor me on positive solutions of equations....can any of you two help me? please???

OpenStudy (saifoo.khan):

i think u typed the wrong ques

OpenStudy (anonymous):

well i dont even think that was a possible equation that even existed...lol. i think ur right

OpenStudy (anonymous):

saifoo is right if it is 5x^2 the answer would be 9/5 and 0

OpenStudy (saifoo.khan):

hmm. :)

OpenStudy (anonymous):

hmmmm *twicee. :-))

OpenStudy (anonymous):

interesting. well can any of you two give me an equation and explain to me how you get the positive solution of it?

OpenStudy (anonymous):

factor out an x x(5x-9)=0

OpenStudy (anonymous):

Yeah!!!! A problem like that is what I have on my paper. Now...how do I get the positive solution of this?

OpenStudy (anonymous):

then set x=0 and 5x-9=0

OpenStudy (anonymous):

Yeah. Then what?

OpenStudy (anonymous):

will we know that x=0 is one answer. then let's solve 5x-9=0. once we solve that we get 5x=9 then divide 5 by both sides we get 9/5

OpenStudy (anonymous):

dose that make sence?

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

can we do another problem if you don't mind?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i did this one but i want to make sure it's right: 4x^2 - 36 = 0

OpenStudy (anonymous):

i got: the positive solution is 3.

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

that is what I got... for 4x^2-36=0

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

sorry that is what I got

OpenStudy (anonymous):

No, it's 9/5. You factor out an x, and you use the zero product rule, setting x = 0 and (5x-9)=0, since, algebraically, you can do that. So, solving for x, you get x = 0 and x = 9/5

OpenStudy (anonymous):

LOL. but i was talking about a different equation.....4x squared - 36 = 0

OpenStudy (anonymous):

mknight94 we arewwe are doing a different problem

OpenStudy (anonymous):

*not the 1st one

OpenStudy (anonymous):

smartchickk59 you are right. are you sure that it is not 4x^2-36x=0?

OpenStudy (anonymous):

not with the x after the 36

OpenStudy (anonymous):

no its: 4x^2-36=0

OpenStudy (anonymous):

3 is the answer

OpenStudy (anonymous):

Did you need any more help?

OpenStudy (anonymous):

Oh no thank you...not for now. But thank you so muchh for your help and your patince....!!! :) Thanks.

OpenStudy (anonymous):

*patience